Tyler has two savings accounts that his grandparents opened for him. The two accounts pay 10% and 12% in annual interest; there is $700 more in the account that pays 12% than there is in the other account. If the total interest for a year is $260, how much money does he have in each account?
Answer by richwmiller(17219)

We don't know the total invested.
We know the difference in the two accounts.
We know that the interest for the two accounts is $260
0.12*x+0.1*y=260
We know that the account at 12% has $700 more.
x=700+y
We substitute for x
0.12*(700+y)+0.1*y=260
We multiply out
84+0.12y+0.1*y=260
We combine like terms.
0.22*y=176
Isolate y
y=$800 at 10%
x=700+y
Calculate x
x=$1500 at 12%
Now,we know the total invested is: 2300
Total invested $1500+$800=$2300
We check
0.12*1500+0.1*800=260
180+80=260
260=260
Since this statement is TRUE and neither x nor y is negative all is well.